CALCULUS(MAXIMA AND MINIMA): A rectangular trough is 3m wide and 15m long.water flows in at the rate of 3m^3/min.how fast is the surface rising?
*show the illustration

Answer by Alan3354(69443)

A rectangular trough is 3m wide and 15m long. water flows in at the rate of 3m^3/min. how fast is the surface rising?
===========
The cross-section is constant --> rate of rise is constant.
---
Vol = L*W*H = 15*3*H = 45H
dV/dt = 45*(dH/dt) = 3
dH/dt = 1/15 m/minute
--------
Not really a calculus problem.